OpenStudy (anonymous):

An experiment consists of selecting a card at random from a 52 card deck. Find the probability of the event that a spade or a queen is drawn. Give the answer as a fraction.

OpenStudy (anonymous):

OR means ADD And means MULTIPLY so in this case you add the probabilities There are 4 queens in a deck so that is = 4/52 There are 13 spades = 13/52 can you do it from there ?

OpenStudy (anonymous):

once you have the answer though, do you put the two together?

OpenStudy (anonymous):

4/52 + 13/52 = 17/52

OpenStudy (anonymous):

is that an answer choice ?

OpenStudy (anonymous):

I wish this was multiple choice :\